# When “A Helpful Assistant” Is Not Really Helpful: Personas in System Prompts Do Not Improve Performances of Large Language Models

### v2

@@ 1 Introduction
In this study, we aim to answer four major research questions: (1) Does adding personas to system prompts help improve model performance on objective tasks? (2) Does the social construct of the persona affect model performance? (3) What factors could potentially explain the effect of personas on model performance? (4) Can we automatically identify the best roles for prompting? Through our analysis, we find that, in general, prompting with personas has no or small negative effects on model performance compared with the control setting where no persona is added. This result is consistent across four popular LLMs, suggesting that adding personas into system prompts may not help improve the model’s performance. To further understand the relative differences among personas, we analyze the social attributes of personas, including role type, gender, and domain alignment. We find that gender-neutral, in-domain, and work-related roles lead to better performance than other types of roles, but with relatively small effect sizes, suggesting that the social construct of the persona may not fully explain the consequential performance differences.
To understand the potential mechanisms behind the relative performance differences caused by different personas, we further analyze the word frequency of the persona, the perplexity, and the similarity of the prompt-questions pairs. Overall, we observe that personas with high-frequency words lead to relatively better model performance. Furthermore, while the similarity between the persona and the question is the strongest predictor of final performance, the correlation between prompt-question similarity and prediction accuracy remains low. Overall, our results suggest that word frequency, perplexity, and prompt-question similarity do not explain much of the performance gaps between different personas. To further uncover the potential mechanism behind persona-based prompting, we try to approach this from another perspective: automatic persona search. The hypothesis is that if the effect of persona varies in a systematic way, we should be able to identify which persona might result in better performance automatically. We test a list of automatic persona selection strategies ranging from simple random selection to fine-tuning RoBERTa-based classifiers. We find that the effect of each persona on model performance varies across questions, presenting a challenge in reliably identifying which persona would consistently yield better performance. Such a result suggests that the effect of personas on LLMs’ performance might largely be random.
@@ 2 Related work
Personas are fundamental in human society and day-to-day interactions (Heiss, 2017; Goffman, 2016). personas define the norm of human interactions and affect human behaviors in various contexts (Sunstein, 1996). Two prominent types of personas are interpersonal roles which are deeply embedded in interpersonal relationships (Berscheid, 1994) (e.g., mother and friend), and professional/occupational roles that fulfill certain social functions or provide certain services in the society (e.g., driver and teacher) (Bucher and Strauss, 1961; Brante, 1988). As suggested by Wolfensberger (2000), “People largely perceive themselves and each other in terms of their roles.” Given the importance of personas in human interactions and recent advances in persona-based agents (Wang et al., 2023; Pataranutaporn et al., 2021; Gupta et al., 2023), understanding LLMs’ role-playing capabilities and the effect of personas hold significance to both the NLP community and the general public. Previous literature examines the influence of in-context impersonation in vision-based reasoning tasks Salewski et al. (2024) and shows that providing sociodemographic information to models is beneficial for subjective NLP tasks in zero-shot settings, though the extent of these benefits varies depending on model types, sizes, datasets, and prompt formats Beck et al. (2024).
@@ 3 Experiment Setting
The overall goal of our study is to test whether adding personas in prompts affects LLMs’ performances. In this section, we describe the experiment setup to examine the persona effect on models’ performances.
@@ 3.1 Dataset
We use a subset of MMLU Hendrycks et al. (2021) for all of our experiments. MMLU is a dataset designed for multitask language understanding and has been widely used as an essential benchmark for evaluating LLMs. It features multiple-choice questions that probe knowledge across a diverse set of subjects, ranging from natural sciences and social sciences to business and law. We choose MMLU as our test dataset because (1) it has been widely used for benchmarking LLMs, (2) it includes questions from diverse disciplines, enabling us to test the effect of prompting with domain-aligned personas, and (3) questions across different domains follow similar formats, which reduces potential confounds.
@@ 3.3 Persona
To excessively evaluate the effect of personas on model performance, we carefully curate a large and diverse list of personas that are actively used in people’s daily interactions. We first collect over 300 personas based on several existing studies (Garg et al., 2018; Massey et al., 2015; Choi et al., 2021), WordNet (Miller, 1995), and our own ad-hoc social role list. We manually examine the roles to remove uncommon roles that are rarely used in daily life, such as “ganger” as a hyponym for “boss”. Our final social role set includes 162 personas, of which 112 roles are occupations, and the remaining are interpersonal relationship roles. Table 4 in the Appendix shows the full list of roles in our experiment.
@@ 3.4 Models
4 Does Prompting with Personas Improve LLMs’ Performance?
To assess whether adding personas helps improve model performance in answering factual questions, we fit a mixed-effects regression model that uses the added persona to predict the inference accuracy, controlling a random effect for each model. The control setting, where no role is added to the system prompts, is used as the reference category. Figure 2 shows the first and last 10 coefficients ranked by magnitude. The coefficients for all roles are detailed in Section B in the Appendix. We observe no significant differences between the best-performing personas and the control setting. Additionally, linear regression results for each model are also listed in Appendix B.
@@ 5 Are Certain Personas Better Than Others?
While adding a persona might not be better than the control setting where no role is added, in practice, LLM service providers or users may still need to define the role of the system for various reasons (e.g., security and language styles). Therefore, it is still worth testing whether different categories of personas could lead to systematically different performances.
@@
For interpersonal roles, we analyze 16 aligned roles and categorize them as masculine, feminine, or neutral, resulting in seven masculine roles, seven feminine roles, and two gender-neutral roles. Table 5 in the Appendix shows the mapping of gender and roles. Such a setting allows us to control the effects of role types and reveal the nuanced effects of gender. We employ a mixed-effects model to analyze the relationship between accuracy and gender, with “accuracy” as the dependent variable, “gender” as an independent categorical variable of values “masculine”, “feminine” and “neutral”, and we include a random effect for each model to account for potential variability across different models. As shown in Figure 6, gender-neutral roles perform significantly better than gendered roles and masculine roles perform slightly worse than feminine roles with a small effect size.
@@
For occupational roles, we use the percentages of workers belonging to each gender in 65 occupational roles, extracted from historical US census data Garg et al. (2018). We fit a similar mixed-effects model with the percentage of masculine workers as the independent variable and include random intercepts for each model. The coefficient of “Masculine”, the percentage of masculine workers for each occupation, is -0.000579. The associated p-value is 0.561, indicating that the gender percentage is not a significant predictor of model performance. The results of the two mixed-effects models for gender impact collectively lead to the conclusion that the gender nature of personas has a very limited impact on the models’ performance in terms of accuracy.
@@
Our experiment includes 162 roles from 7 groups: work, school, social, family, romantic, occupation, and AI. These categories distinguish roles based on the social relationships and settings they typically involve. Does the effect of personas on LLMs’ performance vary with the role categories? To answer this question, we run a mixed effect model using the personas’ role category to predict the prediction accuracy for each question, controlling the model as a random effect. We found that work- and work-related roles are associated with better performances than other types of roles, especially AI and occupational roles. However, the effect sizes are relatively small, suggesting that the category of persona does not have a large effect on LLMs’ performances.
@@
6 Why Certain Personas Lead to Higher Accuracies?
Why do certain personas lead to better performance than others? Despite the complexity across personas, we assess several potential mechanisms. In this section, we propose a method to calculate persona embedding that enables an overall performance comparison. Furthermore, we test whether specific characteristics of the prompt and personas might be driving the behavior: the n-gram frequency of role words, the perplexity of the context prompts, and the similarity between context prompts and questions.
Word Frequency of Personas
For each role, we obtain its n-gram frequency for the period between 2018 and 2019 (the most recent data available) from the Google Ngram Viewer  [^3] https://books.google.com/ngrams/. Frequencies are retrieved for the full role n-gram (e.g., the frequency of “software engineer” as a bigram). Figure 8(a) illustrates the aggregated relationship between accuracy and role word frequency for each model, where each point represents a role and is characterized by its role category. Roles’ n-gram frequencies are weakly correlated to their accuracy for all models, as evidenced by the Pearson correlation coefficients, with the highest correlation of 0.17 for Mistral. This trend suggests that word frequency does not fully explain the effect of personas on model performance.
Prompt-Question Similarity
All of our prompts include two parts: context (e.g., You are talking to your boss) and questions (e.g., Where is the capital of the United States?). Are context prompts that are semantically more similar to the questions more likely to generate accurate answers? To answer this question, we utilize MiniLM (Wang et al., 2020) from Sentence-BERT package (Reimers and Gurevych, 2019) to encode a set of context prompts and full questions with options, and then compute the cosine similarity between the two vectors as a measure of distance between the question and prompt. As shown in Figure 8(b), we observe a weak correlation between similarity and accuracy at the role level. Specifically, the highest correlation is 0.39 on Qwen2.5-7B, whereas the correlation for Qwen2.5-72B is 0.02, suggesting that the effect of similarity is not consistent across different model families and sizes.
Prompt Perplexity
@@ 6 Why Certain Personas Lead to Higher Accuracies?
To perform a comprehensive analysis of all the attributes of roles mentioned previously, we fit a mixed-effects model using three independent variables: the role’s n-gram frequency, prompt-question perplexity, and prompt-question similarity. Random intercepts are included for each model. Table 9 in the Appendix details the regression results. We found that higher frequency, lower perplexity, and higher similarity are generally associated with higher prediction accuracy ($p<0.05$), reflecting similar patterns as shown in the correlation analysis.
@@ 7 Finding the Best Personas for Prompting
In previous sections, we demonstrate that there might not exist a single persona that consistently improves the performance of diverse sets of questions. However, we also observe that personas might help in cases where their domains are aligned with the questions or when they have higher similarities. A natural question arises: instead of using the same role for all questions, could we automatically find the best role for prompting in a specific setting? We experiment with a set of methods to select the best role for prompting LLMs.For the four LLM families used in our analysis, we choose the model with the most parameters for automatic role search experiments.
@@ 8 Conclusion
Incorporating personas into prompts has been an important approach for designing system prompts and enabling role-playing with LLMs. However, the effects of adding these personas on model performance were previously unclear. In this study, we present a systematic analysis of 162 personas in 26 categories to explore how prompting with personas affects model performance. Our analysis shows that, compared with the control setting, adding a persona does not necessarily improve an LLM’s performance on objective tasks. On the contrary, it might actually hurt the models’ overall performance in some situations. Furthermore, while for individual questions, there exists a specific persona that can lead to the correct answer, and aggregating these personas can result in large performance gains, identifying the best role remains challenging, with most selection strategies performing similarly to random selection. Such a result suggests that the effect of personas on model performance can be largely unpredictable. Overall, our study provides a new computational pipeline to evaluate the effect of adding personas on LLMs’ performances. Our findings can help inform the future design of system prompts and role-playing strategies with LLMs. All data, results, and experiment code are available at https://github.com/Jiaxin-Pei/Prompting-with-Social-Roles.
@@ Acknowledgments
This work was funded in part by LG AI Research and by the National Science Foundation under Grant No. IIS-2143529. The authors thank Honglak Lee for feedback on early versions of this work.
@@ 9 Limitations
Our study has the following limitations: First, we only studied four open-source LLMs and didn’t include closed-source models like GPT3.5 and GPT4. This is due to the computational cost of running such a large experiment. We will release the script to run the experiment and we welcome other researchers to explore how role-playing affects LLM performance on other models. Second, while we aimed to be comprehensive when selecting the personas, we were not able to experiment with all the personas beyond the 162 ones in our current experiment. We will release the full list of our personas to support future research in this direction. Third, given the computational costs of our experiments, we only used MMLU as our testbed, overlooking other factual question datasets and open-ended questions. While we believe that our current analysis provides important findings regarding how personas affect the models’ performance, we acknowledge this limitation and plan to extend our analysis to more settings.
@@ 10 Ethical Considerations
Our study has the following ethical implications. First, to ensure the robustness of our results, we experimented with 162 roles, 4 prompt templates, and 9 LLMs over 2410 MMLU questions. Running such an experiment is computationally expensive and is likely to result in a substantial release of carbon dioxide. Second, some of our analyses may reinforce existing stereotypes regarding personas. For example, our results suggest that masculine roles lead to better performance than feminine roles, which might inadvertently reinforce traditional gender stereotypes. However, our results also show that gender-neutral roles lead to higher performance than gendered roles, suggesting that developers should consider using gender-neutral roles when creating system prompts. On the other hand, our results also reveal potential model biases originating from implicit societal stereotypes regarding gender roles. We call for future research in this direction to study de-biasing technologies when training or aligning LLMs.
@@ Appendix A Experiment Settings
The dataset and models used in this study, along with their licenses, are listed in Table 2. All of them are open-source, and our use is consistent with their intended purpose. The mapping between sampled subsets of MMLU and their domains are illustrated in Table 3.

### v3

@@ 1 Introduction
In this study, we aim to answer four major research questions: (1) Does adding personas to system prompts help improve model performance on objective tasks? (2) Does the social construct of the persona affect model performance? (3) What factors could potentially explain the effect of personas on model performance? (4) Can we automatically identify the best roles for prompting? Through our analysis, we find that, in general, prompting with personas has no or small negative effects on model performance compared with the control setting where no persona is added. This result is consistent across four popular LLM families, suggesting that adding personas into system prompts may not help improve the model’s performance. To further understand the relative differences among personas, we analyze the social attributes of personas, including role type, gender, and domain alignment. We find that gender-neutral, in-domain, and work-related roles lead to better performance than other types of roles, but with relatively small effect sizes, suggesting that the social construct of the persona may not fully explain the consequential performance differences.
To understand the potential mechanisms behind the relative performance differences caused by different personas, we further analyze the word frequency of the personas, the similarity between prompt-questions pairs, and the perplexity. We observe that personas with high-frequency words lead to relatively better model performance. Furthermore, while the similarity between the persona and the question is the strongest predictor of final performance, the correlation between prompt-question similarity and prediction accuracy remains low. Our results suggest that word frequency, prompt-question similarity, and perplexity do not explain much of the performance gaps between different personas. To further uncover the potential mechanism behind persona-based prompting, we try to approach this from another perspective: automatic persona search. The hypothesis is that if the effect of persona varies in a systematic way, we should be able to identify which persona might result in better performance automatically. We test a list of automatic persona selection strategies, from simple random selection to fine-tuning RoBERTa-based classifiers. The effect of each persona on model performance varies across questions, making it difficult to reliably identify which persona consistently yields better performance. This suggests that the effect of personas on LLMs’ performance might largely be random.
@@ 2 Related work
Personas are fundamental in human society and day-to-day interactions (Heiss, 2017; Goffman, 2016). personas define the norm of human interactions and affect human behaviors in various contexts (Sunstein, 1996). Two prominent types of personas are interpersonal roles which are deeply embedded in interpersonal relationships (Berscheid, 1994) (e.g., mother and friend), and professional/occupational roles that fulfill certain social functions or provide certain services in the society (e.g., driver and teacher) (Bucher and Strauss, 1961; Brante, 1988). As suggested by Wolfensberger (2000), “People largely perceive themselves and each other in terms of their roles.” Given the importance of personas in human interactions and recent advances in persona-based agents (Wang et al., 2023; Pataranutaporn et al., 2021; Gupta et al., 2023), understanding LLMs’ role-playing capabilities and the effect of personas hold significance to both the NLP community and the general public. Previous literature examines the influence of in-context impersonation in vision-based reasoning tasks Salewski et al. (2024) and shows that providing sociodemographic information to models may benefit subjective NLP tasks in zero-shot settings. However, some other studies also point out the potential bias and limitations of such persona and social-demographic-based prompting (Sun et al., 2023; Hu and Collier, 2024; Beck et al., 2024). While existing studies on subjective tasks uncover important patterns in models’ performance changes when prompted with personas, teasing out the effect of personas on models’ performances is hard because of the natural subjectivity of the task. Therefore, our study focuses on objective tasks where the models’ performance change is solely affected by the added persona.
@@ 3 Experiment Setting
Our study aims to test whether adding personas in prompts affects LLMs’ performances. In this section, we describe the experiment setup to examine the persona effect on models’ performances.
@@ 3.1 Dataset
We use a subset of MMLU Hendrycks et al. (2021) for all of our experiments. MMLU, a dataset designed for multitask language understanding, is widely used to benchmark LLMs. It features multiple-choice questions that probe knowledge across a diverse set of subjects, ranging from natural sciences and social sciences to business and law. We choose MMLU as our test dataset because (1) it has been widely used for benchmarking LLMs, (2) it includes questions from diverse disciplines, enabling us to test the effect of prompting with domain-aligned personas, and (3) questions across different domains follow similar formats, which reduces potential confounds.
@@ 3.3 Persona
To excessively evaluate the effect of personas on model performance, we curate a large and diverse list of personas that are actively used in people’s daily interactions. We first collect over 300 personas based on several existing studies (Garg et al., 2018; Massey et al., 2015; Choi et al., 2021), WordNet (Miller, 1995), and our own ad-hoc social role list. We manually examine the roles to remove uncommon roles that are rarely used in daily life, such as “ganger” as a hyponym for “boss”. Our final social role set includes 162 personas, of which 112 roles are occupations, and the remaining are interpersonal relationship roles. Table 4 in the Appendix shows the full list of roles in our experiment.
@@ 3.4 Models
4 Does Prompting with Personas Improve LLMs’ Performance?
To assess whether adding personas helps improve model performance in answering factual questions, we fit a mixed-effects regression model that uses the added persona to predict the inference accuracy, controlling a random effect for each model to account for potential variability across different models. The control setting, where no role is added to the system prompts, is used as the reference category. Figure 2 shows the first and last 10 coefficients ranked by their effect sizes on the models’ performance change compared with the control setting. The coefficients for all roles are detailed in Section B in the Appendix. We observe no significant differences between the best-performing personas and the control setting. Additionally, linear regression results for each model are also listed in Appendix B.
@@ 5 Are Certain Personas Better Than Others?
While adding a persona might not outperform the control setting with no role, in practice, LLM service providers or users may still need to define the role of the system for various reasons (e.g., security and language styles). Therefore, it is still worth testing whether different categories of personas could lead to systematically different performances.
@@
For interpersonal roles, we analyze 16 aligned roles and categorize them as masculine, feminine, or neutral, resulting in seven masculine roles, seven feminine roles, and two gender-neutral roles. Table 5 in the Appendix shows the mapping of gender and roles. Such a setting allows us to control the effects of role types and reveal the nuanced effects of gender. We employ a mixed-effects model to analyze the relationship between accuracy and gender, with “accuracy” as the dependent variable, “gender” as an independent categorical variable of values “masculine”, “feminine” and “neutral”, and we include a random effect for each model. As shown in Figure 6, gender-neutral roles perform significantly better than gendered roles and masculine roles perform slightly better than feminine roles with a small effect size.
For occupational roles, we use the percentages of workers belonging to each gender in 65 occupational roles, extracted from historical US census data Garg et al. (2018). We fit a similar mixed-effects model with the percentage of masculine workers as the independent variable and include random intercepts for each model. The coefficient of “Masculine”, the percentage of masculine workers for each occupation, is -5.79e-4. The associated p-value is 0.561, indicating that the gender percentage is not a significant predictor of model performance. The results of the two mixed-effects models for gender impact collectively lead to the conclusion that the gender nature of personas has a very limited impact on the models’ performance in terms of accuracy.
@@
Our experiment includes 162 roles from 7 groups: work, school, social, family, romantic, occupation, and AI. These categories distinguish roles based on the social relationships and settings they typically involve. Does the effect of personas on LLMs’ performance vary with the role categories? To answer this question, we run a mixed effect model using the personas’ role category to predict the prediction accuracy for each question, controlling the model as a random effect. We found that work- and school-related roles are associated with better performances than other types of roles, especially AI and occupational roles. However, the effect sizes are relatively small, suggesting that the category of persona does not have a large effect on LLMs’ performances.
@@
6 Why Certain Personas Lead to Higher Accuracies?
Why do certain personas lead to better performance than others? Despite the complexity across personas, we assess several potential mechanisms. In this section, we propose a method to calculate persona embedding that enables an overall performance comparison. Furthermore, we test whether specific characteristics of the prompt and personas might be driving the behavior: the n-gram frequency of role words, the similarity between context prompts and questions, and the perplexity of the context prompts.
Word Frequency of Personas
For each role, we obtain its n-gram frequency for the period between 2018 and 2019 (the most recent data available) from the Google Ngram Viewer  [^3] https://books.google.com/ngrams/. Frequencies are retrieved for the full role n-gram (e.g., the frequency of “software engineer” as a bigram). Figure 8(a) illustrates the aggregated relationship between accuracy and role word frequency for each model, where each point represents a role and is characterized by its role category. The n-gram frequencies of roles are weakly correlated with their accuracy across all models, as indicated by the Pearson correlation coefficients. The largest correlation in absolute value is -0.23 for Qwen2.5-7B, and the smallest is -0.01 for Llama3-8B. This trend suggests that word frequency does not fully explain the effect of personas on model performance.
Prompt-Question Similarity
All of our prompts include two parts: context (e.g., You are talking to your boss) and questions (e.g., Where is the capital of the United States?). Are context prompts that are semantically more similar to the questions more likely to generate accurate answers? To answer this question, we utilize MiniLM (Wang et al., 2020) from Sentence-BERT package (Reimers and Gurevych, 2019) to encode a set of context prompts and full questions with options, and then compute the cosine similarity between the two vectors as a measure of distance between the question and prompt. As shown in Figure 8(b), we observe a weak correlation between similarity and accuracy at the role level. Furthermore, the effect of similarity is inconsistent across different model families and sizes. Specifically, the highest correlation is 0.39 on Qwen2.5-7B, whereas the smallest absolute correlation is -0.02 for Qwen2.5-72B.
Prompt Perplexity
@@ 6 Why Certain Personas Lead to Higher Accuracies?
To perform a comprehensive analysis of all the attributes of roles mentioned previously, we fit a mixed-effects model using three independent variables: the role’s n-gram frequency, prompt-question similarity, and prompt-question perplexity. Random intercepts are included for each model. Table 9 in the Appendix details the regression results. We find that higher frequency, higher similarity, and lower perplexity are generally associated with higher prediction accuracy ($p<0.05$), reflecting similar patterns as shown in the correlation analysis.
@@ 7 Finding the Best Personas for Prompting
In previous sections, we demonstrate that there might not exist a single persona that consistently improves the performance of diverse sets of questions. However, we also observe that personas might help in cases where their domains are aligned with the questions or when they have higher similarities. A natural question arises: instead of using the same role for all questions, could we automatically find the best role for prompting in a specific setting? We experiment with a set of methods to select the best role for prompting LLMs. For the four LLM collections used in our analysis, we choose the model with the largest number of parameters for automatic role search experiments.
@@ 8 Conclusion
Incorporating personas into prompts has been an important approach for designing system prompts and enabling role-playing with LLMs. However, the effects of adding these personas on model performance were previously unclear. In this study, we present a systematic analysis of 162 personas in 26 categories to explore how prompting with personas affects model performance. Our analysis shows that, compared with the control setting, adding a persona does not necessarily improve an LLM’s performance on objective tasks. On the contrary, it might actually hurt the models’ overall performance in some situations. Furthermore, while a specific persona may lead to the correct answer for individual questions, and aggregating these personas can result in significant performance gains, identifying the best role remains challenging, as most selection strategies perform similarly to random selection. This result suggests that the effect of personas on model performance can be largely unpredictable. Our study introduces a new computational pipeline to evaluate the impact of adding personas on LLM performance. These findings can help inform the future design of system prompts and role-playing strategies with LLMs. All data, results, and experiment code are available at https://github.com/Jiaxin-Pei/Prompting-with-Social-Roles.
@@ Acknowledgments
This work was funded in part by LG AI Research and by the National Science Foundation under Grant No. IIS-2143529. The authors thank Honglak Lee for his feedback on early versions of this work.
@@ 9 Limitations
Our study has the following limitations: First, we only studied four open-source LLM families and didn’t include closed-source models like GPT3.5 and GPT4. This is due to the computational cost of running such a large experiment. We will release the script to run the experiment and we welcome other researchers to explore how role-playing affects LLM performance on other models. Second, while we aimed to be comprehensive in selecting personas, we were unable to experiment with all possible personas beyond the 162 used in our current experiment. We will release the full list of personas to support future research in this area. Third, due to the computational costs of our experiments, we used only MMLU as our testbed, overlooking other factual question datasets and open-ended questions. While we believe our current analysis offers important insights into how personas affect model performance, we acknowledge this limitation and plan to extend our analysis to additional settings.
@@ 10 Ethical Considerations
Our study has the following ethical implications. First, to ensure the robustness of our results, we experimented with 162 roles, 4 prompt templates, and 9 LLMs across 2,410 MMLU questions. Running such experiments is computationally expensive and likely contributes to a substantial carbon dioxide footprint. Second, some of our analyses may reinforce existing stereotypes regarding personas. For example, our results suggest that masculine roles lead to better performance than feminine roles, which might inadvertently reinforce traditional gender stereotypes. However, our results indicate that gender-neutral roles result in higher performance than gendered roles, suggesting that developers should prioritize gender-neutral roles when creating system prompts. Additionally, our results reveal potential model biases stemming from implicit societal stereotypes regarding gender roles. We call for future research to study de-biasing technologies when training or aligning LLMs.
@@ Appendix A Experiment Settings
The dataset and models used in this study, along with their licenses, are listed in Table 2. All of them are open-source, and our use aligns with their intended purpose. The mapping between sampled MMLU subsets and their domains is shown in Table 3.